Circular No. 6410

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ae IN NGC 5775 S. Benetti, European Southern Observatory (ESO); and T. Zwitter,

University of Ljubljana, communicate: “A low-signal-to-noise CCD spectrogram (range 332–905 nm, resolution 0.8 nm) of SN 1996ae (cf. IAUC 6409), obtained on May 27.98 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph), is dominated by an Hα emission (with a complex profile) superimposed on a red continuum. The Hα consists of three components: an unresolved one centered at 660.1 nm, an intermediate one centered at 660.1 nm (FWHM ∼ 2950 km/s), and a broad one centered at 665.4 nm (FWHM ∼ 13 350 km/s). This profile is reminiscent of that shown by the peculiar type-II (IIn) supernova 1988Z (see Turatto et al. 1993, MNRAS 262, 128), ∼ 100 days after maximum, but with the substantial difference that the broadest component of Hα emission in the SN 1996ae spectrum shows an unusual redward displacement with respect to the narrower components. Hβ is present in emission (centered at 489.1 nm), showing only the intermediate component (FWHM ∼ 2050 km/s). A weak He I 587.6-nm emission (FWHM ∼ 2000 km/s) is streaked by a very intense Na I D interstellar band (from which a redshift of 1850 km/s is derived for NGC 5775).”

SUPERNOVA 1996ab IN ANONYMOUS GALAXY G. V. Williams, Harvard-Smithsonian Center for Astrophysics, pro-

vides his measurement of SN 1996ab from a CCD image obtained on May 26.221 UT by J. McEnery with the Mt. Hopkins 1.2-m reflector: α = 15^{h}21^{m}08^{s}.35, δ = +27^{o}55′36′′.8 (equinox 2000.0), which is 9′′.3 west and 15′′.6 north of the galaxy’s center. Williams notes that the semiaccurate position for SN 1996ab on IAUC 6405 actually refers to the galaxy’s center.

GRO J1655–40 R. Hunstead and D. Campbell-Wilson, University of Sydney, report on

the return of radio emission from GRO J1655–40, accompanying the recent increase in x-ray (IAUC 6393) and optical/ultraviolet (IAUC 6406) flux: “A 12-hr observation with the Molonglo Observatory Synthesis Telescope (MOST) at 843 MHz on May 28.60 UT (mid-exposure) showed a point source of 55 ± 5 mJy at the optical position. No source was present at this position to a limit of 3 mJy in an earlier MOST image on May 10.”

1996 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green